OpenStudy (linn99123):

Solve for the variable. a - 16 = -10

OpenStudy (midhun.madhu1987):

a - 16 = -10 a = -10 + 16 a = ?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

its six

OpenStudy (anonymous):

i added it up bc ur supposed to change the - sign to a plus sign

OpenStudy (gorv):

add 16 on both side

OpenStudy (anonymous):

no u dont do that

OpenStudy (anonymous):

y would u add 16 on both sides gorv

OpenStudy (pottersheep):

a - 16 = -10 a = -10 + 16 a = 6. Remember that signs switch when you move across the equal sign

OpenStudy (anonymous):

u not no mathlete gorv

OpenStudy (anonymous):

pottersheep im glad ur on my side